# BOPIM: Bayesian Optimization for influence maximization on temporal networks

### v4

@@ 2.5 Discussion
$\mathsf{BOPIM}$ is much faster than a standard greedy approach since we only need to obtain $N_{0}+B$ realizations of the costly objective function, as opposed to $O(nk)$ times for a greedy algorithm. Thus, the number of evaluations of the objective function for $\mathsf{BOPIM}$ is independent of $k$. Moreover, it is trivial to adapt the algorithm to other diffusion mechanisms. In practice, the diffusion model should be chosen with great care based on the application at hand. But since the diffusion model only appears in the evaluation of the objective function, and the objective function is treated as a “black-box”, changing the diffusion model has no impact on the implementation of the method. This differs from, e.g., Dynamic Degree (Murata and Koga,, 2018), which only works for the SI model, or Erkol et al., (2020), which is based on the SIR model.
@@ 3.1 Data
The following experiments are conducted on four real-world networks. Reality ($n=64$, $m=722$) (Eagle and Pentland,, 2006), Hospital ($n=75$, $m=1139$) (Vanhems et al.,, 2013), Bluetooth ($n=136$, $m=5949$) (Kotz and Henderson,, 2004) and Conference 2 ($n=199$, $m=1550$) (Chaintreau et al.,, 2007) are all proximity networks where an edge exists if two people are close to each other. In order to discretize the continuous appearing and disappearing of edges, each network is aggregated into $T$ temporal snapshots of equal duration. We note that these networks are relatively small due to the slow speed of the greedy algorithm. Therefore, in the Supplemental Materials, we also conduct an experiment using the proposed method on the DNC email network (Rossi and Ahmed,, 2015) with $n=1891$ nodes to demonstrate its scalability. Additionally in the Supplemental Materials, we also quantify the fit of the surrogate model to the true objective function.
@@ 3.2 Settings
We simulate the influence diffusion process and select seed sets for IM using $\mathsf{BOPIM}$. The main metrics of comparison are the proportion of nodes influenced at the conclusion of the spreading process, as well as the total algorithmic run time. We consider the proposed algorithm with both Hamming and Jaccard kernel, denoted $\mathsf{BOPIM-H}$ and $\mathsf{BOPIM-J}$, respectively, with $N_{0}=5$ and $B=20$. A greedy algorithm is also used, considered the “gold standard” in IM. Erkol et al., (2022) proved that the SI model of diffusion is monotone and sub-modular so we use the CELF (Leskovec et al.,, 2007) step to increase the speed of the greedy algorithm. We also compare with random sampling seed nodes (Random), randomly sampling proportional to a nodes’ degree (Random Degree) and Dynamic Degree (Murata and Koga,, 2018).
All experiments are conducted under the ex post assumption where the future evolution of the network is known when selecting seed nodes as in Murata and Koga, (2018); Osawa and Murata, (2015); Aggarwal et al., (2012); Erkol et al., (2020). While this assumption may be unrealistic in some cases, the main goal of this work is to propose a BO scheme for IM rather than tackling the ex ante problem (Yanchenko et al.,, 2023). We leave this as an important avenue of future work. Finally, to compare the different methods, we vary the number of seed nodes $(k)$, granularity of snapshots $(T)$, and infection parameter $(\lambda)$. Results are averaged over 25 Monte Carlo (MC) replications (save Greedy and Dynamic Degree), and the influence spread results are reported with error bars.
@@ 4.1.1 Modified mean function
Specifically, let the mean function $\mu(\cdot)$ be parameterized by $\bm{\beta}$. Motivated by Baptista and Poloczek, (2018), we assume that the surrogate model is linear in its parameters, i.e.,
@@
Since $n$ is large for most real-world networks, the dimension of $\bm{\beta}$ will also be large. Additionally, it is unlikely that each component of $\bm{\beta}$ is significant, i.e., many nodes inclusion in the seed set does not lead to significant influence spread. Thus, we endow $\bm{\beta}$ with a sparsity-inducing shrinkage prior. As in Baptista and Poloczek, (2018), we use the Horseshoe (HS) prior (Carvalho et al.,, 2009, 2010) which is designed to shrink insignificant coefficient estimates towards zero while still allowing for accurate estimation of the important coefficients. Specifically, the HS prior handles sparsity via a half-Cauchy prior distribution on the variance of the coefficients, i.e.,
@@
$\eta_{j}^{2},\tau^{2}$ $\sim\mathsf{C}^{+}(0,1)$ (8)
where $j=1,\dots,n$ and $\mathsf{C}^{+}(0,1)$ is the standard half-Cauchy distribution. Thus, $\tau^{2}$ represents the global shrinkage and $\eta_{j}^{2}$ controls the local shrinkage of each coefficient. For the Gibbs sampler, we leverage an auxiliary variable formulation (Makalic and Schmidt,, 2015). Please see the Supplemental Materials for complete MCMC details. We also considered the R2D2 prior (Zhang et al.,, 2022; Yanchenko et al.,, 2025) and Dirichlet-Laplace (Bhattacharya et al.,, 2015), but the results were similar, so we focus on the HS prior.
@@ 5 Conclusion
Another important area for future work is constructing the kernel function. Indeed, the surprising results of the Hamming distance’s superior performance indicates that the role of the kernel function in this problem is not fully understood. One potential direction comes from Automatic Relevance Detection (ARD) (Wipf and Nagarajan,, 2007; Neal,, 2012). This approach enforces sparsity in the model via the covariance function and has already enjoyed success in BO papers (e.g., Papenmeier et al.,, 2023). Finally, considering the ex ante IM task is another important open-problem.
@@ 1. Proof of Hamming Kernel being positive semi-definite
Thus, the covariance matrix is a scaled Gram matrix, implying positive $s$quare
@@ 2. COMBO algorithm
While, technically, $\mathsf{COMBO}$ only requires constructing the combinatorial graph, all results in Oh et al., (2019) directly rely on constructing this graph using the graph Cartesian product. Since we cannot construct the graph in this way for our specific problem, we are unable to use $\mathsf{COMBO}$ for the IM task. We consider it an interesting avenue of future work to extend the ideas of $\mathsf{COMBO}$ to constrained input spaces.
@@ 4. Connection with Trust regions
The greedy algorithm can be seen as a relative of the Trust Region (TR) methodology, made popular in e.g., Eriksson et al., (2019); Wan et al., (2021); Papenmeier et al., (2023). Let $\bm{x}^{*}$ be the current best solution, i.e., seed set which has yielded the largest influence spread. Then $\mbox{TR}_{L}(\bm{x}^{*})$ is the Trust region of radius $L$ around $\bm{x}^{*}$, defined as all points within Hamming distance $L$ of $\bm{x}^{*}$, i.e.,
@@ 5. IM for large networks
To demonstrate the performance of $\mathsf{BOPIM}$ on large networks, we present the results for IM on the DNC email network (Rossi and Ahmed,, 2015) with $n=1891$ nodes. We fix $\lambda=0.05$ and $T=10$ while varying $k=25,50,\dots,100$. Because of the long run times, we only consider $\mathsf{BOPIM-H}$ using 5 MC repetitions, while still using 25 for Random and Random Degree. The influence spread results and run times are in Figure 9. We can see that the proposed method provides comparable results to the greedy algorithm, but is significantly faster. While $\mathsf{BOPIM-H}$ performs similarly to Random Degree, it greatly outperforms Dynamic Degree.
@@ 7. Initial sampling ablation study
In this section, we consider an alternative sampling scheme for the initial observations in the BO algorithm. We remark that nodes with high degrees on the temporally aggregated network, $\tilde{G}=\cup_{t}G_{t}$, are good candidates for the optimal seed set. Additionally, we would like to sample nodes that are “far apart” on the graph to ensure better influence spread. Therefore, instead of randomly sampling $\bm{x}_{i}$ from $\mathcal{X}$, we sample nodes proportional to a metric which accounts for degree and distance from the other nodes in the seed set. For the first node, we sample it proportional to the degree on the temporally aggregated network. Mathematically, this means sampling from a distribution with probability mass function $\mathsf{P}(Z=j)=d_{j}/\sum_{k}d_{k}$ where $d_{j}$ is the degree of node $j$ on $\tilde{G}$ for $Z\in\{1,\dots,n\}$. For the remaining $K-1$ nodes, we compute the average distance between the node $i$, and nodes already sampled for the seed set, call it $\nu_{i}^{(S)}$. We standardize both of these measures by converting to a $z$-score, i.e.,
@@ 9. Horseshoe prior Gibbs sampler
In this section, we report the Horseshoe prior (Polson and Scott,, 2010) Gibbs sampler, using the auxiliary variable formulation of Makalic and Schmidt, (2015).
@@
    Sample $\bm{\beta}|\bm{\lambda},\bm{\nu},\tau^{2},\sigma^{2},{\bf Y}\sim\mathsf{N}(\bm{\mu},\sigma^{2}{\bf V})$ where $\bm{\mu}={\bf V}{\bf X}^{T}{\bf Y}$, ${\bf V}=({\bf X}^{T}{\bf K}^{-1}{\bf X}+{\bf S}^{-1})^{-1}$,
@@
    Sample $\sigma^{2}|\bm{\beta},\bm{\eta},\bm{\nu},\tau^{2},{\bf Y}\sim\mathsf{IG}(a_{1}+(N_{0}+n)/2,b_{1}+(({\bf Y}-{\bf X}\bm{\beta})^{T}{\bf K}^{-1}({\bf Y}-{\bf X}\bm{\beta})+\bm{\beta}^{T}{\bf S}^{-1}\bm{\beta})/2)$.

### v5

@@ 2.5 Discussion
$\mathsf{BOPIM}$ is much faster than a standard greedy approach since we only need to obtain $N_{0}+B$ realizations of the costly objective function, as opposed to $O(nk)$ times for a greedy algorithm. Thus, the number of evaluations of the objective function for $\mathsf{BOPIM}$ is independent of $k$. Moreover, it is trivial to adapt the algorithm to other diffusion mechanisms. In practice, the diffusion model should be chosen with great care based on the application at hand. But since the diffusion model only appears in the evaluation of the objective function, and the objective function is treated as a “black-box”, changing the diffusion model has no impact on the implementation of the method. This differs from, e.g., Dynamic Degree (Murata and Koga, 2018), which only works for the SI model, or Erkol et al. (2020), which is based on the SIR model.
@@ 3.1 Data
The following experiments are conducted on four real-world networks. Reality ($n=64$, $m=722$) (Eagle and Pentland, 2006), Hospital ($n=75$, $m=1139$) (Vanhems et al., 2013), Bluetooth ($n=136$, $m=5949$) (Kotz and Henderson, 2004) and Conference 2 ($n=199$, $m=1550$) (Chaintreau et al., 2007) are all proximity networks where an edge exists if two people are close to each other. In order to discretize the continuous appearing and disappearing of edges, each network is aggregated into $T$ temporal snapshots of equal duration. We note that these networks are relatively small due to the slow speed of the greedy algorithm. Therefore, in the Supplemental Materials, we also conduct an experiment using the proposed method on the DNC email network (Rossi and Ahmed, 2015) with $n=1891$ nodes to demonstrate its scalability. Additionally in the Supplemental Materials, we also quantify the fit of the surrogate model to the true objective function.
@@ 3.2 Settings
We simulate the influence diffusion process and select seed sets for IM using $\mathsf{BOPIM}$. The main metrics of comparison are the proportion of nodes influenced at the conclusion of the spreading process, as well as the total algorithmic run time. We consider the proposed algorithm with both Hamming and Jaccard kernel, denoted $\mathsf{BOPIM-H}$ and $\mathsf{BOPIM-J}$, respectively, with $N_{0}=5$ and $B=20$. A greedy algorithm is also used, considered the “gold standard” in IM. Erkol et al. (2022) proved that the SI model of diffusion is monotone and sub-modular so we use the CELF (Leskovec et al., 2007) step to increase the speed of the greedy algorithm. We also compare with random sampling seed nodes (Random), randomly sampling proportional to a nodes’ degree (Random Degree) and Dynamic Degree (Murata and Koga, 2018).
All experiments are conducted under the ex post assumption where the future evolution of the network is known when selecting seed nodes as in Murata and Koga (2018); Osawa and Murata (2015); Aggarwal et al. (2012); Erkol et al. (2020). While this assumption may be unrealistic in some cases, the main goal of this work is to propose a BO scheme for IM rather than tackling the ex ante problem (Yanchenko et al., 2023). We leave this as an important avenue of future work. Finally, to compare the different methods, we vary the number of seed nodes $(k)$, granularity of snapshots $(T)$, and infection parameter $(\lambda)$. Results are averaged over 25 Monte Carlo (MC) replications (save Greedy and Dynamic Degree), and the influence spread results are reported with error bars.
@@ 4.1.1 Modified mean function
Specifically, let the mean function $\mu(\cdot)$ be parameterized by $\bm{\beta}$. Motivated by Baptista and Poloczek (2018), we assume that the surrogate model is linear in its parameters, i.e.,
@@
Since $n$ is large for most real-world networks, the dimension of $\bm{\beta}$ will also be large. Additionally, it is unlikely that each component of $\bm{\beta}$ is significant, i.e., many nodes inclusion in the seed set does not lead to significant influence spread. Thus, we endow $\bm{\beta}$ with a sparsity-inducing shrinkage prior. As in Baptista and Poloczek (2018), we use the Horseshoe (HS) prior (Carvalho et al., 2009, 2010) which is designed to shrink insignificant coefficient estimates towards zero while still allowing for accurate estimation of the important coefficients. Specifically, the HS prior handles sparsity via a half-Cauchy prior distribution on the variance of the coefficients, i.e.,
@@
$\eta_{j}^{2},\tau^{2}$ $\sim\mathsf{C}^{+}(0,1)\$(8)
where $j=1,\dots,n$ and $\mathsf{C}^{+}(0,1)$ is the standard half-Cauchy distribution. Thus, $\tau^{2}$ represents the global shrinkage and $\eta_{j}^{2}$ controls the local shrinkage of each coefficient. For the Gibbs sampler, we leverage an auxiliary variable formulation (Makalic and Schmidt, 2015). Please see the Supplemental Materials for complete MCMC details. We also considered the R2D2 prior (Zhang et al., 2022; Yanchenko et al., 2025) and Dirichlet-Laplace (Bhattacharya et al., 2015), but the results were similar, so we focus on the HS prior.
@@ 5 Conclusion
Another important area for future work is constructing the kernel function. Indeed, the surprising results of the Hamming distance’s superior performance indicates that the role of the kernel function in this problem is not fully understood. One potential direction comes from Automatic Relevance Detection (ARD) (Wipf and Nagarajan, 2007; Neal, 2012). This approach enforces sparsity in the model via the covariance function and has already enjoyed success in BO papers (e.g., Papenmeier et al., 2023). Finally, considering the ex ante IM task is another important open-problem.
@@ 1. Proof of Hamming Kernel being positive semi-definite
Thus, the covariance matrix is a scaled Gram matrix, implying positive semi-definiteness. $\square$
@@ 2. COMBO algorithm
While, technically, $\mathsf{COMBO}$ only requires constructing the combinatorial graph, all results in Oh et al. (2019) directly rely on constructing this graph using the graph Cartesian product. Since we cannot construct the graph in this way for our specific problem, we are unable to use $\mathsf{COMBO}$ for the IM task. We consider it an interesting avenue of future work to extend the ideas of $\mathsf{COMBO}$ to constrained input spaces.
@@ 4. Connection with Trust regions
The greedy algorithm can be seen as a relative of the Trust Region (TR) methodology, made popular in e.g., Eriksson et al. (2019); Wan et al. (2021); Papenmeier et al. (2023). Let $\bm{x}^{*}$ be the current best solution, i.e., seed set which has yielded the largest influence spread. Then $\mbox{TR}_{L}(\bm{x}^{*})$ is the Trust region of radius $L$ around $\bm{x}^{*}$, defined as all points within Hamming distance $L$ of $\bm{x}^{*}$, i.e.,
@@ 5. IM for large networks
To demonstrate the performance of $\mathsf{BOPIM}$ on large networks, we present the results for IM on the DNC email network (Rossi and Ahmed, 2015) with $n=1891$ nodes. We fix $\lambda=0.05$ and $T=10$ while varying $k=25,50,\dots,100$. Because of the long run times, we only consider $\mathsf{BOPIM-H}$ using 5 MC repetitions, while still using 25 for Random and Random Degree. The influence spread results and run times are in Figure 9. We can see that the proposed method provides comparable results to the greedy algorithm, but is significantly faster. While $\mathsf{BOPIM-H}$ performs similarly to Random Degree, it greatly outperforms Dynamic Degree.
@@ 7. Initial sampling ablation study
In this section, we consider an alternative sampling scheme for the initial observations in the BO algorithm. We remark that nodes with high degrees on the temporally aggregated network, $\tilde{G}=\cup_{t}G_{t}$, are good candidates for the optimal seed set. Additionally, we would like to sample nodes that are “far apart” on the graph to ensure better influence spread. Therefore, instead of randomly sampling $\bm{x}_{i}$ from $\mathcal{X}$, we sample nodes proportional to a metric which accounts for degree and distance from the other nodes in the seed set. For the first node, we sample it proportional to the degree on the temporally aggregated network. Mathematically, this means sampling from a distribution with probability mass function $\mathsf{P}(Z=j)=d_{j}/\sum_{k}d_{k}$ where $d_{j}$ is the degree of node $j$ on $\tilde{G}$ for $Z\in\{1,\dots,n\}$. For the remaining K1K-1 nodes, we compute the average distance between the node $i$, and nodes already sampled for the seed set, call it $\nu_{i}^{(S)}$. We standardize both of these measures by converting to a $z$-score, i.e.,
@@ 9. Horseshoe prior Gibbs sampler
In this section, we report the Horseshoe prior (Polson and Scott, 2010) Gibbs sampler, using the auxiliary variable formulation of Makalic and Schmidt (2015).
@@
    Sample $\bm{\beta}|\bm{\lambda},\bm{\nu},\tau^{2},\sigma^{2},{\bf Y}\sim\mathsf{N}(\bm{\mu},\sigma^{2}{\bf V})$ where $\bm{\mu}={\bf V}{\bf X}^{T}{\bf Y}$, ${\bf V}=({\bf X}^{T}(\gamma{\bf K}+{\bf I})^{-1}{\bf X}+{\bf S}^{-1})^{-1}$,
@@
    Sample $\sigma^{2}|\bm{\beta},\bm{\eta},\bm{\nu},\tau^{2},{\bf Y}\sim\mathsf{IG}(a_{1}+(N_{0}+n)/2,b_{1}+(({\bf Y}-{\bf X}\bm{\beta})^{T}(\gamma{\bf K}+{\bf I})^{-1}({\bf Y}-{\bf X}\bm{\beta})+\bm{\beta}^{T}{\bf S}^{-1}\bm{\beta})/2)$.